A Complete Bibliography of Publications in Supercomputing Frontiers and Innovations

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

11 November 2017 Version 1.00

Title word cross-reference

3 [MKYZ16, SJMLJ17].

2 [LPL^{+}15]. 2.0 [CHC^{+}16]. 2014 [RNL15].

Accelerated [DHK^{+}14]. Acceleration [CGBS^{+}15, KM15, YY17]. Access [vWEH^{+}16]. acoustic [MKYZ16]. across [CHC^{+}16]. adaptive [DK14]. Agent [SJMLJ17]. Agent-Based [SJMLJ17]. Aggregates [NPM15]. agnostic [FOL^{+}14, MHW^{+}16]. Algebra [DAA^{+}15]. Algebraic [YTK14]. Algorithm [AFK^{+}16, MKYZ16]. Algorithmic [VAD15, MOD15]. Algorithms [CDP^{+}15]. AlgoWiki [AFK^{+}16, VAD15]. Analysis [LPL^{+}15, SG16, TKOM17, Moh14]. Analytical [KKL14]. Analyzing [Kun16]. Application [JVT16, NS16, YY17, vWEH^{+}16]. Applications [MBC^{+}15, SWAB14, CGBS^{+}15, GRS16]. Approach [Gus16, HE17, VMC^{+}14, ACM^{+}14, EDDZ16]. Approaches [KLAJ16]. Architecture [BS17, FOL^{+}14]. Architectures [AOML^{+}15, LCHC15, VMC^{+}14]. Arithmetic [GY17]. art [Moh14]. Artificial [SK16]. Asian [SPJ17]. Assessment [LCHC15]. Astrophysics [GKS^{+}14]. Autonomic [HPC^{+}15]. Aware [AOML^{+}15, VMC^{+}14, KBG16].

Backup [Gua16]. Based [SJMLJ17, EDDZ16, KM15, KNI17]. Basis [JVT16]. Beating [GY17]. become [RNL15]. beyond [BS17]. Big [ACM^{+}14, MST^{+}14, BAL17]. body [NS16]. Borderless [BCH^{+}15]. Bridge [LCHC15].

calculation [NS16]. Case [KM15, NPM15, KLAJ16]. Cell [MBS^{+}16]. Center [NPM15]. CFD [YY17]. Challenge [BAL17]. Challenges [BBB^{+}15, BMM^{+}15, Moh14]. Chebfun [EDDZ16]. Cholesky [AFK^{+}16]. Climate [KKL16, FOL^{+}14]. Cloud [BCH^{+}15, LCHC15]. Cluster [JVT16, RNL15]. clusters [Las14]. Co [GKS^{+}14]. Co-design [GKS^{+}14]. Collective [HM14]. Combinatorial [KLAJ16]. Communication [HM14, KBG16, YTK14]. Complexity [YTK14]. Compression [KKL16, SCH^{+}14]. Computation [Gus16]. Computations [AOML^{+}15]. computer [LDFK16]. Computing [AABC15, BCH^{+}15, BMM^{+}15, CHC^{+}16, Gua16, KLY^{+}17, SPJ17, SCH^{+}14, SAB17, Las14, MBC^{+}15, SKAB14]. concept [NDL^{+}17]. Congestion [MVL^{+}16]. Consumption [SWAB14]. continents [CHC^{+}16]. continual [NS16]. Convergence [BAL17]. Cooling [MDS^{+}16, LDFK16]. Core [Jen17a, KLAJ16]. coupled [ACM^{+}14]. CPU [MBS^{+}16, SG16, UWK^{+}16]. Create [SPJ17]. Creating [MOD15]. Critical [MLC15]. CSC [RNL15]. CUDA [NS16]. cultural [EDDZ16].

D [MKYZ16, SJMLJ17]. Data [ACM^{+}14, BCH^{+}15, Chi15, EDDZ16, Gua16, KNI17, KKL16, Kun16, MST^{+}14, SCH^{+}14, SG16, MHW^{+}16, BAL17]. data-structure [MHW^{+}16]. Database [SPJ17]. Databases [WD16]. Decomposition [AFK^{+}16, JVT16]. Defined [Gua16]. Dense [AOML^{+}15, DAA^{+}15]. Design [HE17, KLY^{+}17, GKS^{+}14]. Development [BDG^{+}17, VO17]. Different [TKOM17]. direct [DK14]. Discovery [TWEL17]. Distance [AOML^{+}15]. Distance-Aware [AOML^{+}15]. Distribute [SPJ17]. distributing [CHC^{+}16]. Do [MDS^{+}16]. Domain [JVT16]. Driven [DHK^{+}14]. Dynamic [MVL^{+}16].

E-scale [Gua16]. Early [DK14]. Easy [vWEH^{+}16]. EBD [MST^{+}14]. edge [KCK^{+}16, MOD15]. Effects [CHHBS17]. Efficiency [BBB^{+}15, MDS^{+}16, RNL15]. Efficient [KM15, CDP^{+}15]. Embedded [KM15]. Encyclopedia [VAD15]. Energy [AABC15, BBB^{+}15, CDP^{+}15, HM14, KM15, MDS^{+}16, SWAB14, TWEL17]. Energy-Efficient [KM15, CDP^{+}15]. Environment [HPC^{+}15, TKOM17]. equations [MKYZ16]. Era [SCH^{+}14]. Evaluation [KNI17, DK14]. Example [Jen17a]. Exascale [CGG^{+}14, Chi15, DFG^{+}15, HPC^{+}15, KKL14, MLC15, NZS16, SCH^{+}14, SKAB14]. Expenses [KKL14]. Exploration [Chi15, KNI17, UWK^{+}16]. Exploring [CHHBS17]. Extract [WD16]. Extreme [BDG^{+}17, BS17, HE17, TWEL17, MST^{+}14]. Extreme-scale [BDG^{+}17].

Fabric [BCH^{+}15]. Facing [BAL17]. Factorizations [DHK^{+}14]. factors [RGL14]. Fast [YTK14]. Fault [Gua16]. feasibility [KBG16]. Features [VAD15]. few [NS16]. few-body [NS16]. Feynman [NS16]. field [MKYZ16]. File [Kun16]. Fine [BS17]. Fine-Grain [BS17]. First [VMC^{+}14, LDFK16]. Flight [YY17]. Floating [GY17]. Flow [VO17]. Fold [SJMLJ17]. Fonton [BS17]. Foreword [MD15]. Formats [Kun16]. Foundations [BDG^{+}17]. Fourier [JVT16]. FPGA [KM15]. FPGA-based [KM15]. FPGAs [LDFK16]. Framework [KNI17, VO17, DTKV15]. Freely [SPJ17]. fundamental [RGL14]. Fusion [TWEL17]. future [Moh14].

Game [GY17]. Generation [MST^{+}14]. Genome [LPL^{+}15]. Genomic [BCH^{+}15, SPJ17]. Global [BCH^{+}15]. GPU [YY17]. GPUs [JVT16]. Grain [BS17]. Graph [KM15]. graphs [MOD15]. Green500 [RNL15]. greenest [RNL15]. ground [NS16]. GUI [vWEH^{+}16].

Hardware [TKOM17, MHW^{+}16]. hardware-independent [MHW^{+}16]. Harvesting [NPM15]. Haswell [LCHC15]. HCA [KBG16]. heritage [EDDZ16]. Heterogeneous [DAA^{+}15, Las14, LPL^{+}15]. hierarchical [Las14]. High [BCH^{+}15, CHC^{+}16, SPJ17, SAB17]. Higher [MDS^{+}16]. HPC [BAL17, CGBS^{+}15, LCHC15, SK16, SWAB14, TWEL17, vWEH^{+}16]. HPCG [Jen17a]. Human [LPL^{+}15]. Hybrid [MBS^{+}16, Las14].

I/O [SK16]. ICON [VO17]. Illustrated [Kun16]. imaging [EDDZ16, GRS16]. impact [SG16]. Implementation [KLY^{+}17, FOL^{+}14, MBS^{+}16]. Implementing [HM14]. Improve [MDS^{+}16]. In-Situ [KNI17, VO17]. independent [MHW^{+}16]. Inescapable [NZS16]. InfiniBand [DK14, LCHC15, MVL^{+}16]. InfiniCloud [BCH^{+}15, CHC^{+}16]. InfiniCortex [BCH^{+}15, NDL^{+}17]. Inflammation [SJMLJ17]. Infrastructure [MST^{+}14]. INMOST [DTKV15]. Insect [YY17]. Instances [LCHC15]. integral [MKYZ16]. integrals [NS16]. Integration [VO17]. Intel [LCHC15]. interconnect [MOD15]. ISAAC [MHW^{+}16].

Kit [BDG^{+}17].

L [RNL15]. L-CSC [RNL15]. Langford [KLAJ16]. Large [KLY^{+}17, NPM15, NZS16, VO17, DK14]. Large-Scale [NZS16, VO17, DK14]. Law [BS17]. Leading [KCK^{+}16]. Leading-edge [KCK^{+}16]. Level [MDS^{+}16]. Leveraging [BCH^{+}15]. Library [KBG16]. Light [NPM15]. Light-Harvesting [NPM15]. Linear [DAA^{+}15]. Liquid [MDS^{+}16, LDFK16]. list [RNL15]. Live [SSG^{+}15]. Local [JVT16].

Machines [DFG^{+}15]. Making [NZS16]. Management [MVL^{+}16]. Many [KLAJ16]. Many-Core [KLAJ16]. Matrix [AOML^{+}15, Jen17a]. Measurement [AABC15, Moh14]. mechanisms [CGBS^{+}15]. Medicine [SPJ17]. Memory [HM14, KYG17, MS14]. merging [EDDZ16]. Message [Jen17b]. method [MKYZ16, MBS^{+}16, NS16, YTK14]. Methods [GKS^{+}14]. MOD [MOD15]. Model [AFK^{+}16, DHK^{+}14, SKAB14]. Model-Driven [DHK^{+}14]. modeling [DTKV15, MKYZ16]. Models [DAA^{+}15, Jen17a, FOL^{+}14]. Module [Jen17a]. Molecular [NPM15]. Monitoring [SG16]. monochromatic [MKYZ16]. Moore [BS17]. MPI [CGBS^{+}15, Gua16, KBG16]. Multi [NPM15]. Multi-Scale [NPM15]. Multicore [DHK^{+}14]. Multipole [YTK14].

Necessary [SPJ17]. Neo [LPL^{+}15]. Neo-heterogeneous [LPL^{+}15]. Nesus [BBB^{+}15, BMM^{+}15]. Networks [MVL^{+}16, SK16, DK14]. Neural [SK16]. Next [MST^{+}14]. Non [Gua16]. Non-stop [Gua16]. November [RNL15]. NR [Gua16]. NR-MPI [Gua16]. nuclei [NS16]. NUMA [AOML^{+}15]. Numbers [Gus16]. Numerical [GKS^{+}14, DTKV15].

O [SK16]. Observable [NZS16]. offloading [KBG16]. One [DHK^{+}14]. One-Sided [DHK^{+}14]. Open [VAD15]. OpenStack [BCH^{+}15]. Operations [HM14]. Opportunities [BMM^{+}15, MS14]. Optimization [LPL^{+}15]. Optimizing [Jen17a, RNL15]. Own [GY17].

Paradigms [DFG^{+}15]. Parallel [AFK^{+}16, DTKV15, DAA^{+}15, GKS^{+}14, KBG16, MKYZ16, VAD15, Las14, Moh14]. Parallelized [LPL^{+}15]. Particle [KNI17, UWK^{+}16, MBS^{+}16]. Particle-in-Cell [MBS^{+}16]. Patterns [HE17]. PDE [Jen17a]. Year [MST^{+}14]. Performance [BCH^{+}15, BS17, CHHBS17, CHC^{+}16, HPC^{+}15, KNI17, LCHC15, MLC15, SK16, SPJ17, SG16, SAB17, TWEL17, TKOM17, FOL^{+}14, Moh14, RGL14, SKAB14]. Phenomena [VO17]. Phi [MBS^{+}16]. Photosynthetic [NPM15]. Physics [GKS^{+}14]. Pipeline [LPL^{+}15]. Plasma [GKS^{+}14, MBS^{+}16]. platform [DTKV15]. platforms [Las14]. Point [GY17]. Population [SPJ17]. Portability [TWEL17]. Portable [MLC15, FOL^{+}14]. Posit [GY17]. possible [SG16]. Power [SWAB14, RNL15]. Precision [SPJ17]. Predicting [SK16, SWAB14]. prediction [RGL14]. Preparing [KCK^{+}16]. Problem [KLAJ16]. Problems [KM15, KLAJ16, MS14]. Processing [AFK^{+}16, KYG17, KCK^{+}16, TKOM17]. Processing-in-Memory [KYG17]. Processing-in-Storage [KYG17]. Production [NDL^{+}17, WD16]. Programmer [Gua16]. Programming [DFG^{+}15, DAA^{+}15, KLY^{+}17, LPL^{+}15, SSG^{+}15]. Project [AFK^{+}16, BBB^{+}15, BMM^{+}15]. Proof [NDL^{+}17]. Proof-of-concept [NDL^{+}17]. Properties [Kun16, SG16]. Protocol [Jen17b]. PULSAR [KLY^{+}17].

Radical [Gus16]. Ray [UWK^{+}16]. Re [LPL^{+}15]. Re-sequencing [LPL^{+}15]. Real [Gus16]. Receive [Jen17b]. Reconfigurable [LDFK16]. removal [MOD15]. Rendering [KNI17]. Repair [SJMLJ17]. Require [DFG^{+}15]. requirements [KBG16]. Research [MS14]. Resilience [BMM^{+}15, CGG^{+}14, HE17]. Resilient [Gua16]. Resources [vWEH^{+}16]. Restore [Gua16]. routing [DK14]. Runtime [HM14, KNI17, SAB17, VMC^{+}14]. Runtime-Aware [VMC^{+}14]. Runtimes [DFG^{+}15].

Sampling [Kun16]. Sandy [LCHC15]. Scalability [RGL14]. Scalable [Moh14]. Scale [HE17, KLY^{+}17, NPM15, NZS16, TWEL17, VO17, BDG^{+}17, DK14, Gua16]. Scaling [SWAB14]. Scheduling [CHHBS17]. Science [BAL17, TWEL17]. Scientific [ACM^{+}14, BDG^{+}17, Kun16, SSG^{+}15]. Seismic [TKOM17]. sequencing [LPL^{+}15]. Server [MDS^{+}16]. Sided [DHK^{+}14]. Simulation [JVT16, SJMLJ17, SSG^{+}15, YY17, MBS^{+}16]. Simulations [TWEL17, UWK^{+}16]. Simultac [BS17]. Simultaneous [Jen17b]. Situ [KNI17, KCK^{+}16, SJMLJ17, UWK^{+}16, VO17, WD16, MHW^{+}16]. SLOWER [SKAB14]. SMOD [MOD15]. SoCs [KM15]. Software [BDG^{+}17, LPL^{+}15, SAB17, TWEL17, DTKV15]. South [SPJ17]. Sparse [Jen17a]. Spectral [JVT16]. STAR [Jen17b]. state [Moh14]. state-of-the-art [Moh14]. states [NS16]. Statistical [Kun16]. Stealing [AOML^{+}15]. steerable [MHW^{+}16]. Step [NZS16]. stop [Gua16]. Storage [KYG17, KKL14]. strategy [FOL^{+}14]. Strong [SWAB14]. structure [MHW^{+}16]. Structured [HE17]. Study [KKL14, NPM15, KBG16]. Super [Gua16]. Supercomputer [GRS16, LPL^{+}15, RNL15]. Supercomputers [KCK^{+}16].

Supercomputing [NPM15]. Supporting [Gua16]. Survey [AABC15, SCH^{+}14, SAB17]. sustainable [CGBS^{+}15]. System [BMM^{+}15, KLY^{+}17, MVL^{+}16, MDS^{+}16]. Systems [BBB^{+}15, CDP^{+}15, DHK^{+}14, DAA^{+}15, Gua16, KKL14, MS14, NZS16, SAB17, LDFK16].

Task [CHHBS17]. TaskInsight [CHHBS17]. Technologies [MST^{+}14, GRS16]. technology [NS16]. Temperatures [MDS^{+}16]. TH [LPL^{+}15]. TH-2 [LPL^{+}15]. their [SG16]. tomographic [GRS16]. Tools [AABC15, ACM^{+}14, Moh14]. topologies [MOD15]. Tracing [UWK^{+}16]. Tradeoffs [HM14]. Transmit [Jen17b]. Trends [BBB^{+}15]. Turbulence [TWEL17].

Ultrascale [AABC15, BBB^{+}15, BMM^{+}15, CDP^{+}15, MBC^{+}15]. Ultrasound [JVT16]. Unsteady [VO17]. Upcoming [KCK^{+}16]. update [CGG^{+}14]. Usage [SG16]. Using [JVT16, SK16, SPJ17, Kun16].

Variants [YTK14]. Visualization [ACM^{+}14, MLC15, SJMLJ17, VO17, WD16, MHW^{+}16]. Vocal [SJMLJ17]. Volume [KNI17].

Weak [SWAB14]. weather [FOL^{+}14]. within [BMM^{+}15]. Work [AOML^{+}15]. Workflows [BAL17]. workstations [Las14]. world [RNL15].

Xeon [MBS^{+}16]. xSDK [BDG^{+}17].

Yottabyte [MST^{+}14]. Yottabyte/Year [MST^{+}14].

References

Almeida:2015:EMT

[AABC15] Francisco Almeida, Javier Arteaga, Vicente Blanco, and Alberto Cabrera. Energy measurement tools for ultrascale computing: A survey. Supercomputing Frontiers and Innovations, 2(2):64–76, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                                   tronic). URL http://superfri.
                                                   org/superfri/article/view/
                                                   45.

Artigues:2014:SBD

[ACM^{+}14] Antoni Artigues, Fernando Martin Cucchietti, Carlos Tripiana Montes, David Vicente, Hadrien Calmet, Guillermo Marin, Guillaume Houzeaux, and Mariano Vazquez. Scientific big data visualization: a coupled tools approach. Supercomputing Frontiers and Innovations, 1(3):4–18, ???? 2014. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                                   URL http://superfri.org/
                                                   superfri/article/view/29.

Antonov:2016:PPM

[AFK^{+}16] Alexander S. Antonov, Alexey V. Frolov, Hiroaki Kobayashi, Igor N. Konshin, Alexey M. Teplov, Vadim V. Voevodin, and Vladimir V. Voevodin. Parallel processing model for Cholesky decomposition algorithm in AlgoWiki Project. Supercomputing Frontiers and Innovations, 3(3):61–70, ???? 2016.

CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
110.

Al-Omairy:2015:DMC

[AOML^{+}15] Rabab Al-Omairy, Guillermo Miranda, Hatem Ltaief, Rosa M. Badia, Xavier Martorell, Jesus Labarta, and David Keyes. Dense matrix computations on NUMA architectures with distance-aware work stealing. Supercomputing Frontiers and Innovations, 2(1):49–72, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
49.

Badia:2017:WSC

[BAL17] Rosa M. Badia, Eduard Ayguade, and Jesus Labarta. Workflows for science: a challenge when facing the convergence of HPC and Big Data. Supercomputing Frontiers and Innovations, 4(1):27–47, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
125.

Bagein:2015:EEU

[BBB^{+}15] Michel Bagein, Jorge Barbosa, Vicente Blanco, Ivona Brandic, Samuel Cremer, Sebastien Fremal, Helen Karatza, Laurent Lefevre, Toni Mastelic, Ariel Oleksiak, Anne-Cecile Orgerie, Georgios L. Stavrinides, and Sebastien Varrette. Energy efficiency for ultrascale systems: Challenges and trends from Nesus Project. Supercomputing Frontiers and Innovations, 2(2): 105–131, ???? 2015. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/48.

Ban:2015:ILG

[BCH^{+}15] Kenneth Hon Kim Ban, Jakub Chrzeszczyk, Andrew Howard, Dongyang Li, and Tin Wee Tan. InfiniCloud: Leveraging the global InfiniCortex fabric and OpenStack cloud for borderless high performance computing of genomic data. Supercomputing Frontiers and Innovations, 2(3):14–27, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
76.

Bartlett:2017:XFT

[BDG^{+}17] Roscoe Bartlett, Irina Demeshko, Todd Gamblin, Glenn Hammond, Michael Allen Heroux, Jeffrey Johnson, Alicia Klinvex, Xiaoye Li, Lois Curfman McInnes, J. David Moulton, Daniel Osei-Kuffuor, Jason Sarich, Barry Smith, James Willenbring, and Ulrike Meier Yang. xSDK foundations: Toward an extreme-scale scientific software development kit. Supercomputing Frontiers and Innovations, 4(1):69–82, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
127.

Bouvry:2015:RWU

[BMM^{+}15] Pascal Bouvry, Rudolf Mayer, Jakub Muszyński, Dana Petcu, Andreas Rauber, Gianluca Tempesti, Tuan Trinh, and Sébastien Varrette. Resilience within ultrascale computing system: Challenges and opportunities from Nesus Project. Supercomputing Frontiers and Innovations, 2(2):46–63, ???? 2015. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/46.

Brodowicz:2017:SFF

[BS17] Maciej Brodowicz and Thomas Sterling. Simultac Fonton: A fine-grain architecture for extreme performance beyond Moore’s Law. Supercomputing Frontiers and Innovations, 4(2):27–37, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/134.

Carretero:2015:EEA

[CDP^{+}15] Jesus Carretero, Salvatore Distefano, Dana Petcu, Daniel Pop, Thomas Rauber, Gudula Rünger, and David E. Singh. Energy-efficient algorithms for ultrascale systems. Supercomputing Frontiers and Innovations, 2(2):77–104, ???? 2015.

CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
41.

Carretero:2015:AMM

[CGBS^{+}15] Jesus Carretero, Javier Garcia-Blas, David E. Singh, Florin Isaila, Alexey Lastovetsky, Thomas Fahringer, Radu Prodan, Peter Zangerl, Christi Symeonidou, Afshin Fassihi, and Horacio Pérez-Sánchez. Acceleration of MPI mechanisms for sustainable HPC applications. Supercomputing Frontiers and Innovations, 2(2):28–45, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
35.

Cappello:2014:TER

[CGG^{+}14] Franck Cappello, Al Geist, William Gropp, Sanjay Kale, Bill Kramer, and Marc Snir. Toward exascale resilience: 2014 update. Supercomputing Frontiers and Innovations, 1(1):5–28, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
14.

Chrzeszczyk:2016:IDH

[CHC^{+}16] Jakub Chrzeszczyk, Andrew Howard, Andrzej Chrzeszczyk, Ben Swift, Peter Davis, Jonathan Low, Tin Wee Tan, and Kenneth Ban. InfiniCloud 2.0: distributing high performance computing across continents. Supercomputing Frontiers and Innovations, 3(2):54–71, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
95.

Ceballos:2017:ESE

[CHHBS17] Germán Ceballos, Andra Hugo, Erik Hagersten, and David Black-Schaffer. Exploring scheduling effects on task performance with TaskInsight. Supercomputing Frontiers and Innovations, 4(3):91–98, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
146.

Childs:2015:DEE

[Chi15] Hank Childs. Data exploration at the exascale. Supercomputing Frontiers and Innovations, 2(3):5–13, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
78.

Dongarra:2015:PPM

[DAA^{+}15] Jack Dongarra, M. Abalenkovs, A. Abdelfattah, M. Gates, A. Haidar, J. Kurzak, P. Luszczek, S. Tomov, I. Yamazaki, and A. YarKhan. Parallel programming models for dense linear algebra on heterogeneous systems. Supercomputing Frontiers and Innovations, 2(4):67–86, ????

2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        90.

DaCosta:2015:EMR

[DFG^{+}15] Georges Da Costa, Thomas Fahringer, Juan Antonio Rico Gallego, Ivan Grasso, Atanas Hristov, Helen D. Karatza, Alexey Lastovetsky, Fabrizio Marozzo, Dana Petcu, Georgios L. Stavrinides, Domenico Talia, Paolo Trunfio, and Hrachya Astsatryan. Exascale machines require new programming paradigms and runtimes. Supercomputing Frontiers and Innovations, 2(2):6–27, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        44.

Dongarra:2014:MDO

[DHK^{+}14] Jack Dongarra, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Asim YarKhan. Model-driven onesided factorizations on multicore accelerated systems. Supercomputing Frontiers and Innovations, 1(1):85–115, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        15.

Daryin:2014:EED

[DK14] Alexander N. Daryin and Anton A. Korzh. Early evaluation of direct large-scale InfiniBand networks with adaptive routing. Supercomputing Frontiers and Innovations, 1(3):56–69, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
33.

Danilov:2015:PSP

[DTKV15] Alexander A. Danilov, Kirill M. Terekhov, Igor N. Konshin, and Yuri V. Vassilevski. Parallel software platform IN-MOST: a framework for numerical modeling. Supercomputing Frontiers and Innovations, 2(4):55–66, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
87.

ElAfrit:2016:DMC

[EDDZ16] Mariem El Afrit, Yann Le Du, Rafaël Del Pino, and Guolin Zhang. Data merging for the cultural heritage imaging based on Chebfun approach. Supercomputing Frontiers and Innovations, 3(3):71–83, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
111.

Fuhrer:2014:TPP

[FOL^{+}14] Oliver Fuhrer, Carlos Osuna, Xavier Lapillonne, Tobias Gysi, Ben Cumming, Mauro Bianco, Andrea Arteaga, and Thomas Christoph Schulthess. Towards a performance portable, architecture agnostic implementation strategy for weather and climate models. Supercomputing Frontiers and Innovations, 1(1):45–62, ???? 2014. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                        URL http://superfri.org/
                                        superfri/article/view/17.

Glinskiy:2014:CDP

[GKS^{+}14] Boris M. Glinskiy, Igor M. Kulikov, Alexey V. Snytnikov, Alexey A. Romanenko, Igor G. Chernykh, and Vitaly A. Vshivkov. Co-design of parallel numerical methods for plasma physics and astrophysics. Supercomputing Frontiers and Innovations, 1(3):88–98, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        26.

Goncharsky:2016:STT

[GRS16] Alexander V. Goncharsky, Sergey Y. Romanov, and Sergey Y. Seryozhnikov. Supercomputer technologies in tomographic imaging applications. Supercomputing Frontiers and Innovations, 3(1):41–66, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        98.

Guang:2016:NMN

[Gua16] Suo Guang. NR-MPI: A nonstop and fault resilient MPI supporting programmer defined data backup and restore for E-scale super computing systems. Supercomputing Frontiers and Innovations, 3(1):4–21, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
89.

Gustafson:2016:RAC

[Gus16] John L. Gustafson. A radical approach to computation with real numbers. Supercomputing Frontiers and Innovations, 3(2):38–53, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
94.

Gustafson:2017:BFP

[GY17] John L. Gustafson and Isaac T.

Yonemoto. Beating floating point at its own game: Posit arithmetic. Supercomputing Frontiers and Innovations, 4 (2):71–86, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/137.

Hukerikar:2017:RDP

[HE17] Saurabh Hukerikar and Christian Engelmann. Resilience design patterns: A structured approach to resilience at extreme scale. Supercomputing Frontiers and Innovations, 4(3):4–42, ????

2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        138.

Hoefler:2014:EMR

[HM14] Torsten Hoefler and Dmitry Moor. Energy, memory, and runtime tradeoffs for implementing collective communication operations. Supercomputing Frontiers and Innovations, 1(2):58–75, ???? 2014. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                        URL http://superfri.org/
                                        superfri/article/view/12.

Huck:2015:APE

[HPC^{+}15] Kevin A. Huck, Allan Porterfield, Nick Chaimov, Hartmut Kaiser, Allen D. Malony, Thomas Sterling, and Rob Fowler. An autonomic performance environment for exascale. Supercomputing Frontiers and Innovations, 2(3):49–66, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        64.

Jennings:2017:CMO

[Jen17a] Earle Jennings. Core module optimizing PDE sparse matrix models with HPCG example. Supercomputing Frontiers and Innovations, 4(2):54–70, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.

org/superfri/article/view/
136.

Jennings:2017:STR

[Jen17b] Earle Jennings. The Simultaneous Transmit And Receive (STAR) Message Protocol. Supercomputing Frontiers and Innovations, 4(2):38–53, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
135.

Jaros:2016:SDD

[JVT16] Jiri Jaros, Filip Vaverka, and Bradley E. Treeby. Spectral domain decomposition using local Fourier basis: Application to ultrasound simulation on a cluster of GPUs. Supercomputing Frontiers and Innovations, 3(3):40–55, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
108.

Kulkarni:2016:HAP

[KBG16] Kedar Kulkarni, Shreeya Badhe, and Geetanjali Gadre. HCA aware parallel communication library: A feasibility study for offloading MPI requirements. Supercomputing Frontiers and Innovations, 3(3):56–60, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/

Kress:2016:PSP

[KCK^{+}16] James Kress, Randy Michael Churchill, Scott Klasky, Mark Kim, Hank Childs, and David Pugmire. Preparing for in situ processing on upcoming leadingedge supercomputers. Supercomputing Frontiers and Innovations, 3(4):49–65, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        115.

Kunkel:2014:ESS

[KKL14] Julian Martin Kunkel, Michael Kuhn, and Thomas Ludwig. Exascale storage systems — an analytical study of expenses. Supercomputing Frontiers and Innovations, 1(1):116–134, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        20.

Kuhn:2016:DCC

[KKL16] Michael Kuhn, Julian Kunkel, and Thomas Ludwig. Data compression for climate data. Supercomputing Frontiers and Innovations, 3(1):75–94, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        101.

Krajecki:2016:MCA

[KLAJ16] Michaël Krajecki, Julien Loiseau, François Alin, and Christophe Jaillet. Many-core approaches to combinatorial problems: case of the Langford Problem. Supercomputing Frontiers and Innovations, 3(2):21–37, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
93.

Kurzak:2017:DIP

[KLY^{+}17] Jakub Kurzak, Piotr Luszczek, Ichitaro Yamazaki, Yves Robert, and Jack Dongarra. Design and implementation of the PUL-SAR programming system for large scale computing. Supercomputing Frontiers and Innovations, 4(1):4–26, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
121.

Kapre:2015:CEF

[KM15] Nachiket Kapre and Pradeep Moorthy. A case for embedded FPGA-based SoCs in energy-efficient acceleration of graph problems. Supercomputing Frontiers and Innovations, 2(3):76–86, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
62.

Kawamura:2017:PER

[KNI17] Takuma Kawamura, Tomoyuki Noda, and Yasuhiro Idomura. Performance evaluation of runtime data exploration framework based on in-situ particle based volume rendering. Supercomputing Frontiers and Innovations, 4(3):43–54, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        142.

Kunkel:2016:ADP

[Kun16] Julian Martin Kunkel. Analyzing data properties using statistical sampling — illustrated on scientific file formats. Supercomputing Frontiers and Innovations, 3(3):34–39, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        106.

Kaplan:2017:PMP

[KYG17] Roman Kaplan, Leonid Yavits, and Ran Ginosar. From processing-in-memory to processingin-storage. Supercomputing Frontiers and Innovations, 4 (3):99–116, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                        URL http://superfri.org/
                                        superfri/article/view/147.

Lastovetsky:2014:HPC

[Las14] Alexey Lastovetsky. Heterogeneous parallel computing: from clusters of workstations to hierarchical hybrid platforms. Supercomputing Frontiers and Innovations, 1(3):70–87, ???? 2014.

CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
32.

Low:2015:PAI

[LCHC15] Jonathan Low, Jakub Chrzeszczyk, Andrew Howard, and Andrzej Chrzeszczyk. Performance assessment of InfiniBand HPC cloud instances on Intel Haswell and Intel Sandy Bridge architectures. Supercomputing Frontiers and Innovations, 2 (3):28–40, ???? 2015. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/74.

Levin:2016:RCS

[LDFK16] Ilya I. Levin, Alexey I. Dordopulo, Alexander M. Fedorov, and Igor A. Kalyaev. Reconfigurable computer systems: from the first FPGAs towards liquid cooling systems. Supercomputing Frontiers and Innovations, 3(1):22–40, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
97.

Liao:2015:NHP

[LPL^{+}15] Xiangke Liao, Shaoliang Peng, Yutong Lu, Yingbo Cui, Chengkun Wu, Heng Wang, and Jiajun Wen. Neo-heterogeneous programming and parallelized optimization of a human genome re-sequencing analysis software pipeline on TH-2 supercomputer. Supercomputing Frontiers and Innovations, 2(1):73–83, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
39.

Mihajlovic:2015:AUC

[MBC^{+}15] Milan Mihajlovic, Lars Ailo Bongo, Raimondas Ciegis, Neki Frasheri, Dragi Kimovski, Peter Kropf, Svetozar Margenov, Maya Neytcheva, Thomas Rauber, Gudula Runger, Roman Trobec, Roel Wuyts, Roman Wyrzykowski, and Jing Gong. Applications for ultrascale computing. Supercomputing Frontiers and Innovations, 2(1):19–48, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
43.

Meyerov:2016:HCX

[MBS^{+}16] Iosif B. Meyerov, Sergey I. Bastrakov, Igor A. Surmin, Alexey V. Bashinov, Evgeny S. Efimenko, Artem V. Korzhimanov, Alexander A. Muraviev, and Arkady A. Gonoskov. Hybrid CPU + Xeon Phi implementation of the Particle-in-Cell method for plasma simulation. Supercomputing Frontiers and Innovations, 3(3):5–10, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/

Michalewicz:2015:F

[MD15] Marek Michalewicz and Yuefan Deng. Foreword. Supercomputing Frontiers and Innovations, 2 (3):4, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

Moskovsky:2016:SLL

[MDS^{+}16] Alexander A. Moskovsky, Egor A.

Druzhinin, Alexey B. Shmelev, Vladimir V. Mironov, and Andrey Semin. Server level liquid cooling: Do higher system temperatures improve energy efficiency? Supercomputing Frontiers and Innovations, 3(1):67–74, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
100.

Matthes:2016:SSH

[MHW^{+}16] Alexander Matthes, Axel Huebl, René Widera, Sebastian Grottel, Stefan Gumhold, and Michael Bussmann. In situ, steerable, hardware-independent and data-structure agnostic visualization with ISAAC. Supercomputing Frontiers and Innovations, 3(4):30–48, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
114.

Malovichko:2016:PAM

[MKYZ16] Mikhail S. Malovichko, Nikolay E. Khokhlov, Nikolay B.

Yavich, and Michael S. Zhdanov. Parallel algorithm for 3D modeling of monochromatic acoustic field by the method of integral equations. Supercomputing Frontiers and Innovations, 3(4):74–78, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
120.

Moreland:2015:VEP

[MLC15] Kenneth Moreland, Matthew Larsen, and Hank Childs. Visualization for exascale: Portable performance is critical. Supercomputing Frontiers and Innovations, 2(3):67–75, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
77.

Michalewicz:2015:CIT

[MOD15] Marek T. Michalewicz, Łukasz P. Orłowski, and Yuefan Deng. Creating interconnect topologies by algorithmic edge removal: MOD and SMOD graphs. Supercomputing Frontiers and Innovations, 2(4):16–47, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
80.

Mohr:2014:SPP

[Moh14] Bernd Mohr. Scalable parallel performance measurement and analysis tools — state-of-the-art and future challenges. Supercomputing Frontiers and Innovations, 1(2):108–123, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
18.

Mutlu:2014:RPO

[MS14] Onur Mutlu and Lavanya Subramanian. Research problems and opportunities in memory systems. Supercomputing Frontiers and Innovations, 1 (3):19–55, ???? 2014. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/30.

Matsuoka:2014:EBD

[MST^{+}14] Satoshi Matsuoka, Hitoshi Sato, Osamu Tatebe, Michihiro Koibuchi, Ikki Fujiwara, Shuji Suzuki, Masanori Kakuta, Takashi Ishida, Yutaka Akiyama, Toyotaro Suzumura, Koji Ueno, Hiroki Kanezashi, and Takemasa Miyoshi. Extreme Big Data (EBD): Next generation big data infrastructure technologies towards yottabyte/year. Supercomputing Frontiers and Innovations, 1(2):89–107, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
24.

Mizero:2016:DCM

[MVL^{+}16] Fabrice Mizero, Malathi Veeraraghavan, Qian Liu, Robert D.

Russell, and John M. Dennis. A dynamic congestion management system for InfiniBand networks. Supercomputing Frontiers and Innovations, 3(2):5–20, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
91.

Noaje:2017:IPC

[NDL^{+}17] Gabriel Noaje, Alan Davis, Jonathan Low, Seng Lim, Geok Lian Tan, Łukasz Orłowski, Dominic Chien, Sing-Wu Liou, Tin Wee Tan, Yves Poppe, Kenneth Ban Hon Kim, Andrew Howard, David Southwell, Jason Gunthorpe, and Marek Michalewicz. InfiniCortex — from proof-of-concept to production. Supercomputing Frontiers and Innovations, 4(2):87–102, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/139.

Nemukhin:2015:MSS

[NPM15] Alexander V. Nemukhin, Igor V. Polyakov, and Alexander I. Moskovsky. Multi-scale supercomputing of large molecular aggregates: A case study of the light-harvesting photosynthetic center. Supercomputing Frontiers and Innovations, 2(4):48–54, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.

org/superfri/article/view/
86.

Naumenko:2016:ACT

[NS16] Mikhail A. Naumenko and Vyacheslav V. Samarin. Application of CUDA technology to calculation of ground states of fewbody nuclei by Feynman’s continual integrals method. Supercomputing Frontiers and Innovations, 3(2):80–95, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
102.

Nikitenko:2016:MLS

[NZS16] Dmitry A. Nikitenko, Sergey A.

Zhumatiy, and Pavel A. Shvets. Making large-scale systems observable — another inescapable step towards exascale. Supercomputing Frontiers and Innovations, 3(2):72–79, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
96.

Rosas:2014:SPF

[RGL14] Claudia Rosas, Judit Giménez, and Jesús Labarta. Scalability prediction for fundamental performance factors. Supercomputing Frontiers and Innovations, 1(2):4–19, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/

Rohr:2015:CCO

[RNL15] David Rohr, Gvozden Neskovic, and Volker Lindenstruth. The L-CSC cluster: Optimizing power efficiency to become the greenest supercomputer in the world in the Green500 list of November 2014. Supercomputing Frontiers and Innovations, 2(3):41–48, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        65.

Sterling:2017:SRS

[SAB17] Thomas Sterling, Matthew Anderson, and Maciej Brodowicz. A survey: Runtime software systems for high performance computing. Supercomputing Frontiers and Innovations, 4(1):48–68, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        126.

Son:2014:DCE

[SCH^{+}14] Seung Woo Son, Zhengzhang Chen, William Hendrix, Ankit Agrawal, Wei keng Liao, and Alok Choudhary. Data compression for the exascale computing era — survey. Supercomputing Frontiers and Innovations, 1(2):76–88, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.

org/superfri/article/view/
13.

Stefanov:2016:ACU

[SG16] Konstantin S. Stefanov and Alexey A. Gradskov. Analysis of CPU usage data properties and their possible impact on performance monitoring. Supercomputing Frontiers and Innovations, 3(4):66–73, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
118.

Seekhao:2017:SVA

[SJMLJ17] Nuttiiya Seekhao, Joseph JaJa, Luc Mongeau, and Nicole Y. K. Li-Jessen. In situ visualization for 3D agent-based vocal fold inflammation and repair simulation. Supercomputing Frontiers and Innovations, 4(3):68–79, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/144.

Schmid:2016:PPH

[SK16] Jan Fabian Schmid and Julian M. Kunkel. Predicting I/O performance in HPC using artificial neural networks. Supercomputing Frontiers and Innovations, 3(3):19–33, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/

Sterling:2014:SPM

[SKAB14] Thomas Sterling, Daniel Kogler, Matthew Anderson, and Maciej Brodowicz. SLOWER: A performance model for exascale computing. Supercomputing Frontiers and Innovations, 1(2):42–57, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        10.

Shah:2017:UHP

[SPJ17] Asmi H. Shah, Jonathan D. Picker, and Saumya S. Jamuar. Using high performance computing to create and freely distribute the South Asian Genomic Database, necessary for precision medicine in this population. Supercomputing Frontiers and Innovations, 4(2):4–12, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        132.

Swift:2015:LPS

[SSG^{+}15] Ben Swift, Andrew Sorensen, Henry Gardner, Peter Davis, and Viktor K. Decyk. Live programming in scientific simulation. Supercomputing Frontiers and Innovations, 2(4):4–15, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        72.

Shoukourian:2014:PEP

[SWAB14] Hayk Shoukourian, Torsten Wilde, Axel Auweter, and Arndt Bode. Predicting the energy and power consumption of strong and weak scaling HPC applications. Supercomputing Frontiers and Innovations, 1(2):20–41, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
9.

Tyutlyaeva:2017:SPP

[TKOM17] Ekaterina Olegovna Tyutlyaeva, Sergey Konyukhov, Igor Odintsov, and Alexander Moskovsky. Seismic processing performance analysis on different hardware environment. Supercomputing Frontiers and Innovations, 4 (3):80–90, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

URL http://superfri.org/
superfri/article/view/145.

Tang:2017:PPH

[TWEL17] William Tang, Bei Wang, Stephane Ethier, and Zhihong Lin. Performance portability of HPC discovery science software: Fusion energy turbulence simulations at extreme scale. Supercomputing Frontiers and Innovations, 4(1):83–97, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
128.

Usher:2016:SEP

[UWK^{+}16] Will Usher, Ingo Wald, Aaron Knoll, Michael Papka, and Valerio Pascucci. In situ exploration of particle simulations with CPU ray tracing. Supercomputing Frontiers and Innovations, 3(4):4–18, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        112.

Voevodin:2015:AOE

[VAD15] Vladimir V. Voevodin, Alexander S. Antonov, and Jack Dongarra. AlgoWiki: an open encyclopedia of parallel algorithmic features. Supercomputing Frontiers and Innovations, 2(1):4–18, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        69.

Valero:2014:RAA

[VMC^{+}14] Mateo Valero, Miquel Moreto, Marc Casas, Eduard Ayguade, and Jesus Labarta. Runtimeaware architectures: A first approach. Supercomputing Frontiers and Innovations, 1(1): 29–44, ???? 2014. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                        URL http://superfri.org/
                                        superfri/article/view/19.

Vetter:2017:DIS

[VO17] Michael Vetter and Stephan Olbrich. Development and integration of an in-situ framework for flow visualization of large-scale, unsteady phenomena in ICON. Supercomputing Frontiers and Innovations, 4(3):55–67, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
143.

vanWaveren:2016:EAH

[vWEH^{+}16] Matthijs van Waveren, Ahmed Seif El Nawasany, Nasr Hassanein, David Moon, Niall O’Byrnes, Alain Clo, Karthikeyan Murugan, and Antonio Arena. Easy access to HPC resources through the application GUI. Supercomputing Frontiers and Innovations, 3(3):11–18, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
104.

Whitlock:2016:SVP

[WD16] Brad Joseph Whitlock and Earl P. N. Duque. In situ visualization and production of extract databases. Supercomputing Frontiers and Innovations, 3(4):19–29, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.
org/superfri/article/view/
113.

Yokota:2014:CCF

[YTK14] Rio Yokota, George Turkiyyah, and David Keyes. Communication complexity of the Fast Multipole Method and its algebraic variants. Supercomputing Frontiers and Innovations, 1(1):63–84, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.
                                        org/superfri/article/view/
                                        22.

Yao:2017:AGA

[YY17] Yang Yao and Khoon-Seng Yeo. An application of GPU acceleration in CFD simulation for insect flight. Supercomputing Frontiers and Innovations, 4 (2):13–26, ???? 2017. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic).

                                        URL http://superfri.org/
                                        superfri/article/view/133.